OpenStudy (anonymous):

Let f(x)= {2x^2-4x x<=5/4 {x-(25/8) 5/4

OpenStudy (anonymous):

would it be 1?

OpenStudy (anonymous):

Yup :) 4x-4|x=5/4=5-4=1